Project

General

Profile

Kaon LT Meetings » mtg_23jan19.txt

Garth Huber, 01/20/2023 03:55 PM

 
1
                 Jan 19/23 PionLT/KaonLT Analysis Meeting Notes
2
                 ----------------------------------------------
3
                           (Notes by GH and SJDK)
4

    
5
                    Today: KaonLT will be discussed first
6

    
7
Please remember to post your slides at:
8
https://redmine.jlab.org/projects/kltexp/wiki/Kaon_LT_Meetings
9

    
10
Present
11
-------
12
Regina - Stephen Kay, Garth Huber, Ali Usman, Alicia Postuma, Nathan Heinrich, 
13
   Love Preet, Vijay Kumar
14
JLab - Dave Gaskell
15
Ohio - Jacob Murphy, Julie Roche
16
CUA - Richard Trotta, Tanja Horn
17
FIU - Pete Markowitz
18

    
19
Richard Updates
20
---------------
21
LT-separation code
22
  - making progress on different parts of code
23
  - cuts scripts DONE
24
     - showed some sample t-cuts
25
     - GH - Why are there -ve -t values?
26
        - Plot isn't finalised yet, just for testing -t cut selection
27
        - No selection on the Lambda peak, PID cuts not optimized
28
     - Diamond cut bug, a value crashes the script
29
        - JM - It's looking for first/last bin above some minimum, may not be
30
          finding it
31
  - calculating yields script in progres
32
  - combine SIMC, Dummy, Data - combine.py DONE
33
  - average_kin, calc_unsep Fortran Code DONE
34
  - LTsep, fitting (C++) IN PROGRESS
35
  - plotting scripts IN PROGRESS
36

    
37
  - SIMC script from Bill was outdated, meeting with Bill tomorrow, get correct
38
    version
39
  - Will add SIMC on top of existing plots
40

    
41
  - Combining Left, Right, Center SHMS settings
42
     - Bill and DG had a discussion, apparently the code adds data L/R/C and
43
       simulation L/R/C and combines to get yields
44
     - this is a surprise, since normally we *compute the Exp/MC ratios* first
45
       for L/R/C, and then *combine the ratios* in an error weighted average to
46
       get the combined phi coverge for a setting
47
     - GH: would rather do the ratios method, as was done for the pion analysis
48
     - DG: if the model properly describes the data, both methods should give
49
       the same results, but it is likely the ratios method will cause the MC
50
       model to converge more quickly to the data
51
     - TH: both methods were checked in Fpi-2, and gave same results, but K+
52
       analysis is new and the MC model will start further from the data
53

    
54
     - Should probably get ratios for L/R/C individually and get an error
55
       weighted average
56
        - Try both?
57

    
58
TO DO:
59
- finalize PID, Lumi, SIMC calcs for Heep
60
   - Need to discuss SIMC EM, PM calcs with DG
61
- Will post flowchart to Github, slides online too
62

    
63

    
64
Ali Updates
65
-----------
66
PID studies and HMS Calor/Cher Efficiencies
67
  - Need to finalise efficiences and PID prior to LT-sep, so a high priority
68
  - Looking at HMS Calorimeter and Cherenkov
69
     - Need to get a very clean e- sample
70
     - Use Cherenkov to get good Calo sample and vice versa
71
  - DG: What data is this?
72
  - AU - High eps, 10.6 GeV production data (Autumn 2018)
73
       Q2=3.0, W=2.32, Phms=6.59, S1Xrate=283kHz, 51uA, Run 4865
74
       Run Plan predicted e-Rate=63kHz, pi-Rate=0.5kHz
75
     - Coincidence data, but look at all HMS events
76

    
77
  - 95.10 +/ 0.01% efficiency for HMS Cherenkov
78
  - 95.86 +/- 0.01% effeciency for HMS calorimeter
79

    
80
  - DG: This is a fairly high HMS momentum setting
81
     - Thinks the gas pressure was set to give a 4GeV/c pion threshold (hence
82
       more NPE), since the e-/pi- rate was expected to be favorable
83
       Pi- passing the Cherenkov cut could explain low calorimeter efficiency
84
  - Try a setting with a reasonable ratio but *below 4 GeV on HMS*  Compare
85

    
86
  - DG: *p112 of Burcu's thesis* discusses the effect of a HMS Cher light leak
87
    in the J/Psi data
88
     - Even with cut at 0 NPE, Burcu saw a 3% inefficiency, which was believed
89
       to be due to FADC deadtime caused by the PMT always firing
90
     - NH: comparison of Ali's NPE plot with Burcu's: Burcu has 0PE peak that
91
       seems to be absent in Ali's plot
92
        - Don't *think* we're seeing this?  Possibly the light leak opened up
93
          after the fall 2018 data were taken? 
94
        - Dave G thinks the light leak was fixed in summer 2019
95
        - This implies the *spring 19 KaonLT and summer 2019 PionLT data* will
96
          be *affected*.  Needs investigation, particularly by Vijay.
97

    
98
  - AU: will also look into trimming edges of calorimeter, as discussed by
99
    PeterB
100
     - DG, TH: this should be in the reconstruction code automatically (as
101
       part of tracked variable fiducial cut), no need to do it twice
102

    
103
TO DO:
104
- check position dep in Calor
105
- look at different rates
106

    
107

    
108
Vijay Updates
109
-------------
110
- working on LT-sep code with RT and Bill, trying to understand scripts
111

    
112
- Need to replay PionLT data again
113
  - GH - Please keep notes on *issues Ali presented* as they need separate
114
    investigation for Summer 2019 data
115

    
116

    
117
Alicia Updates
118
--------------
119
pi+ BSA Analysis (from KaonLT data)
120
  - 10.6 GeV data replayed with new error calculations from AU (dummy and LH2)
121
  - only 10.6 GeV will give good results
122
     - High stats, full phi coverage and good polarisation
123

    
124
  - Double checked ROOT error propagation, seems to be handling them correctly,
125
    treated as independent random errors
126

    
127
  - Peter B suggested not doing a dummy subtraction and modifying the fit
128
  - Doing a full dummy target subtraction, separately for each helicity.  This
129
    way there is no need to make any assumptions on AL BSA.  There are lots of
130
    Dummy target statistics, so this seems the best way of handling this
131
     - still need to correct for Dummy target thickness, over-subtracting right
132
       now
133
     - RT: dummy_target_corr = 4.8579
134

    
135
   - see evidence of K+Lambda leakthrough in MM spectrum, need to adjust HGC,
136
     Aerogel cuts
137
     - HGC cut at >5 seems to let in some kaon background
138
     - DG: not concerned about a little K+ contribution, as it is kinematically
139
       unfavored to get into pi+n peak.  However, agree that it is good to
140
       understand the cuts and their effects
141

    
142
  - Fits to Asymmetry Plots
143
    BSA = Asin(phi) / [ 1+ Bcos(phi)+ Ccos(2phi) ]
144
     - second t-bin (-t=0.23-0.30) looks "overfit", particularly in comparison
145
       to the other t-bins, which have more sinusoidal fits
146
     - Tried some new fitting strategies
147
       1) A varies within dA from 1 parameter fit
148
       2) Full fit with restriction 0<B<A, 0<C<A
149
     - method 1 not effective in reducing "overfit"
150
     - method 2 surprisingly effective, ended up with B,C nearly zero
151
     - Some justification for enforcing A > B(and C) would be nice, theory
152
       paper somewhere?
153

    
154
     - JR: Fit with B/C fairly large doesn't really look that bad, how much does
155
       A actually change here?
156
        - If A doesn't change much, will be fine
157
        - Size of A/B/C and their relative errors are very important
158
        - experience from DVCS analysis is that the simplest (sin) fit
159
          *overestimates our knowledge of A* compared to true knowledge, we ended
160
          up listing both results in our publication
161

    
162
     - t-bins will be re-divided after Ali's analysis is further along, they
163
       are not optimized for statistics per bin yet
164

    
165
     - TH: actually first t-bin also has some similar features (wiggles in
166
       phi-distribution) as bin 2, so this might be some consistent feature of
167
       the data that needs to be looked at
168

    
169
     - JR: In future, can you *show how error bars* for the asymmetry are
170
       calculated? (message Julie)
171

    
172
  - New Plots added
173
     - LT' cross section plotted versus -t
174
     - Polar plot of the asymmetry (t/phi, asymmetry is colour scale)
175

    
176
TO DO:
177
- Need to incorporate different dummy target thickness
178
- optimize PID cuts and cuts for mean t,epsilon calcs
179
- Repeat analysis for other (Q2,W) settings
180
- Run Regge/GPD models for comparison to LT' data
181

    
182
Jacob Updates
183
-------------
184
HMS Optics
185
  - Finished all of the cuts for optics settings!
186
     - over 600 cuts for all optics settings set
187
  - Preliminary calibrations for 5.9 setting next week hopefully, as well as
188
    update on 5.8
189

    
190
Nathan Updates - Lots of plots!
191
--------------
192
PionLT Ref Times update
193
  - Following Carlos Yero' procedure
194
  - Start with ADC Ref Time for HMS/SHMS
195
  - Multiplicity == 1 cut for all plots, while CY used >2 cut as he had more
196
    events at ==2 than NH sees
197

    
198
  - HMS 4 peak structure in all COIN runs, 2 peak structures only in SINGLES
199
    running (ELREAL trigger)
200
  - DG: If trigger is ELREAL, *two legs could take timing*, EL-LO or EL-HI,
201
    would explain the two peaks for each set (COIN, SINGLES)
202
  - DG: normally there are separate timing cuts for COIN and SINGLES events
203
     - NH prefers to place one set of cuts, including both sets of peaks, as
204
       there are very few events between the two sets of peaks
205
    - DG: wide timing cuts may cause issues when looking at other detectors, so
206
      this needs to be carefully looked at
207
    - DG: try to *find a run* where HMS rate is high and *see if it looks* the
208
      same as when HMS rates are low, maybe you can get away with loose cut
209

    
210
  - SHMS timing peak structure
211
    - need to figure out which peak is COIN trigger
212
    - presumably COIN comes first (right peak)
213
    - see shifts in RAW trigger time between 2021,22
214
       - Brad replaced a faulty delay module in June 2022, some small timing
215
       - adjustments (see below)
216
   - NH: will need 2 versions of cuts for SINGLE arm data, depending on whether
217
     ELREAL or 3/4 trigger is used
218
       - similarly, all HMS-3/4 runs will need separate timing cuts
219

    
220
- Brad posted an updated trigger flowchart
221
  - Fall 2018-Summer 2019 Trigger Config  https://redmine.jlab.org/attachments/1731
222
  - Fall 2021-Summer 2022 Trigger Update  https://redmine.jlab.org/attachments/1730
223
  - Fall 2021 oscilloscope traces  https://logbooks.jlab.org/entry/3902530
224
  - Summer 2022 oscilloscope traces  https://logbooks.jlab.org/entry/4004019
225

    
226
TO DO:
227
- need to look at TDC reference times for Hodo and DC
228
   - probably need separate cuts for each Hodo TDC
229
- communicated with Cameron Cotton on NGC calib
230
   - now has his code, but hasn't looked at it yet
231

    
232
- JM: should start JM, NH, MJ regular separate meetings soon
233

    
234
Stephen Updates
235
---------------
236
- will try to do some reorganization of Hall C widi
237
   - will have a template for comment by next week
238

    
239
- also do not forget the Hall C Quarterly Analysis meeting next Thursday
240

    
241
Next Meeting
242
------------
243
Thursday Jan 26 @ 17:00 Eastern/16:00 Regina/14:00 Pacific
244
- GH will not be able to attend the first ~30 minutes, hopes to join late,
245
  depending on Air Canada performance
(156-156/413)